OpenStudy (anonymous):

You make a cone from paper by first cutting out a circle with an 8-in. radius. Then you cut along two radii that form a 90° central angle and remove the quarter−circle wedge. You join the two cuts of the three-quarter circle to form the cone. What is the approximate volume of the cone?

OpenStudy (wolf1728):

Cone volume = PI*r²*h/3 The circle has an 8 inch radius and so the slant height of the cone is 8 inches and the radius of the cone base is 6. (See attached graphic) Height of cone² = 8² - 6² Height = sqrt(28) Height = 5.2915026221 SO, volume = (PI*6*6*5.2915) / 3 volume = 199.485 cubic inches